OpenStudy (anonymous):

What is the average rate of change from x = −3 to x = 5?

OpenStudy (anonymous):

I guess that depends on how long it took to get from -3 to 5

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

choices −1 0 1 8

OpenStudy (anonymous):

You need more information. What is y when x is -3? What is y when x is 5?

OpenStudy (anonymous):

well yes the difference between them is 8, so i'd go with 8

OpenStudy (anonymous):

but the question is asked in a very weird way, because there isn't really a 'rate of change'.

OpenStudy (the_fizicx99):

I think your missing some information. There has to be a place hold value for x = -3 and x = 5

OpenStudy (anonymous):

Well, rate of change is slope, so, if this is a line in on the x-axis slope would be zero. And if there is change in y-coordinates it would be m=(x-x1)/(y-y1)

OpenStudy (the_fizicx99):

^ you don't have your y ._.

OpenStudy (anonymous):

Right??

OpenStudy (the_fizicx99):

No.. >.> pretty sure its not like that

OpenStudy (anonymous):

\[\left| -3+5 \right|=8\]

OpenStudy (anonymous):

I'm sure rate of change is slope. 100% sure.

OpenStudy (anonymous):

yes but you need to know y or t or something for a slope

Parth (parthkohli):

Yup, the rate of change is the slope!

OpenStudy (the_fizicx99):

Exactly what I said ._.

OpenStudy (anonymous):

Right so where is y?

OpenStudy (anonymous):

\[\frac{ \Delta y }{ \Delta x }\]

OpenStudy (anonymous):

No, @JTsoccer, where is y in your question?

OpenStudy (the_fizicx99):

What?? x = −3 TO x = 5 I don't get it, ( ._.) You must have been giving some other kind of information..

OpenStudy (anonymous):

Is there a graph with this question, or coordinates, or values for y or t?

OpenStudy (anonymous):

I'd say slope is zero for that problem since why is not given and it's assumed the line is on the x-axis.

OpenStudy (anonymous):

@JTsoccer

OpenStudy (anonymous):

typo, why = y.. lol

OpenStudy (the_fizicx99):

From what I see it, you get the value of of x = -3 and x = -5 and plug them into f(b) - f(a) /b - a

Parth (parthkohli):

There is no rate of change between two lines.

OpenStudy (the_fizicx99):

Without @JTsoccer were going no where here. Peace ~ :3

OpenStudy (anonymous):

@tHe_FiZiCx99 yeah lol